# Towards Swift Serverless LLM Cold Starts with ParaServe

### v1

@@ 10 Related Work
Cold-start optimizations in serverless model serving. There have been plenty of works on reducing the cold start latency in serverless model serving [60, 61, 62, 31, 27]. For instance, FaaSwapfaaswap and ServerlessLLM [27] utilize locality by caching models in local memory or SSDs, while INFless [61] prewarms instances based on historical request patterns. gCROP [63] introduces a on-demand GPU image restore technique that accelerates GPU runtime initialization. Unlike these works, ParaServe primarily targets reducing model fetching latency by aggregating network bandwidth of GPU servers.
Pipeline parallelism. Pipeline parallelism has been widely used to scale GPU resources for training large models [32, 33, 34, 35, 36, 37]. Recently, researchers have also explored its use during inference [38, 64, 65, 66]. AlpaServe [38] observes that model parallelism improves GPU utilization under bursty workloads and introduces a model placement policy when deploying models. HPipe [64] and PipeEdge [66] apply pipeline parallelism to improve the inference performance for LLMs on edge devices. PipeInfer [65] explores speculative inference through asynchronous pipeline parallelism to reduce inference latency. ParaServe leverages pipeline parallelism to reduce cold start latency in serverless LLM serving and further introduces pipeline consolidation, which dynamically scales works in a pipeline parallelism group to improve performance.
LLM serving optimizations. Many LLM serving optimizations have been proposed to improve serving performance [38, 41, 67, 39, 68, 44, 69, 40, 70, 27]. In the area of request scheduling, Orca [68] introduces iteration-level scheduling to run more requests in parallel, while Llumnix [40] employs runtime scheduling to meet user SLOs. FastServe [41] focuses on average job completion time and introduces a preemptive scheduling policy. For key-value cache management, vLLM [44] leverages the concept of virtual memory in operating systems and InfiniGen [69] optimizes KV block placement to improve inference efficiency. Additionally, to mitigate the interference between prefill and decoding phases, DistServe [39] proposes disaggregated inference, whereas SARATHI [70] adopts chunked prefill. ParaServe specifically focuses on reducing cold start latency in serverless LLM serving. Our techniques are orthogonal to existing inference optimizations and we can seamlessly integrate new optimizations into ParaServe.
@@ 11 Conclusion
This paper presents ParaServe, a serverless LLM serving system designed to reduce cold start latency through parallelized model fetching. ParaServe adopts a two-level hierarchical approach to accelerate cold starts, and consolidates pipeline workers into individual workers to ensure high performance for warm requests. Evaluation results show that ParaServe reduces cold start latency by up to 4.7 compared to baselines and improves TTFT SLO attainment by up to 1.74 under various constraints. By integrating cluster-level and worker-level optimizations, ParaServe offers a scalable and efficient solution to meet user-defined SLOs in serverless LLM serving, particularly for latency-sensitive applications.

### v2

@@ 9 Related Work
Cold-start optimizations in serverless model serving. There have been plenty of works on reducing the cold start latency in serverless model serving [71, 72, 73, 43, 30, 63, 31, 33]. For instance, FaaSwap [43] and ServerlessLLM [30] cache models in local memory or SSDs, while INFless [72] and InstaInfer [42] prewarm instances according to historical request patterns. DeepFlow [33] combines model caching and NPU forking to optimize cold start latencies. BlitzScale [31] and Scale [32] achieve fast autoscaling using high-capacity networks, while Medusa [63] speeds up CUDA graph and KV cache construction through state materialization.
HydraServe addresses the cold-start problem in public clouds. Its advantages are twofold. First, it imposes no requirement on the hotness of models and effectively tackles the cold-start problem without relying on caching or worker cooperation. Second, it reduces the end-to-end runtime preparation overhead without any offline profiling.
LLM serving optimizations. Many LLM serving optimizations have been proposed to improve serving performance [51, 57, 74, 55, 75, 40, 76, 56, 77]. In the area of request scheduling, Orca [75] introduces iteration-level scheduling to run more requests in parallel, while Llumnix [56] employs runtime scheduling to meet user SLOs. For key-value cache management, vLLM [40] leverages the concept of virtual memory in operating systems and InfiniGen [76] optimizes KV block placement to improve inference efficiency. Additionally, DistServe [55] disaggregates prefill and decoding phases to mitigate their performance interference. HydraServe specifically focuses on reducing cold start latency in serverless LLM serving and can easily integrate these optimizations.
Pipeline parallelism. Pipeline parallelism has been widely used to scale GPU resources for training large models [45, 46, 47, 48, 49, 50, 58]. Recently, researchers have also explored its usage during inference [51, 52, 53, 54, 32, 31]. While AlpaServe [51] observes that model parallelism improves GPU utilization under load spikes and develops a model placement policy, HPipe [52] and PipeEdge [54] apply pipeline parallelism to improve the inference performance for LLMs on edge devices. Scale [32] also adopts pipeline parallelism to allow scaled-up workers to start inference earlier. HydraServe utilizes pipeline parallelism in a proactive manner that creates multiple workers even if the model only needs one. It further introduces pipeline consolidation to merge pipeline workers during inference.
@@ 10 Conclusion
This paper presents HydraServe, a serverless LLM serving system designed to reduce cold start latency in public clouds. HydraServe targets the most time-consuming stages of a cold start: model fetching and runtime preparation. To optimize model fetching, HydraServe proactively distributes models across multiple servers, alleviating the burden on any single server. Furthermore, HydraServe incorporates pipeline consolidation that merges workers back into individual endpoints to ensure efficient resource usage and high performance for warm requests. For runtime preparation, HydraServe accelerates the local worker startup by overlapping distinct stages. Evaluation results show that HydraServe reduces cold start latency by 1.7– 4.7 compared to baselines and improves TTFT SLO attainment by 1.43–1.74 under various constraints. By integrating these optimizations, HydraServe offers a robust and efficient solution to meet user-defined SLOs in serverless LLM serving, especially for long-tail models.
Acknowledgments. We sincerely thank our shepherd Arvind Krishnamurthy and the anonymous reviewers for their valuable feedback on this paper. This work was supported in part by the National Key Research and Development Program of China under Grant 2022YFB4500700, the Scientific Research Innovation Capability Support Project for Young Faculty under Grant ZYGXQNJSKYCXNLZCXM-I1, the Fundamental Research Funds for the Central Universities, Peking University, and the National Natural Science Foundation of China under Grant 62172008 and 62325201. Xin Jin is the corresponding author. Chiheng Lou, Sheng Qi, Chao Jin, Xuanzhe Liu, and Xin Jin are also with the Key Laboratory of High Confidence Software Technologies (Peking University), Ministry of Education.
Appendix A TPOT SLO Attainment
We provide the TPOT SLO attainment of systems under different CVs in Figure 16. All systems achieve over 95% TPOT SLO attainment in most scenarios, and more than 90% under all CVs and RPS configurations.